In the short run, per capita personal income (PCPI) in the United States grows approximately linearly. In 2009 PCPI was 38.6, and in 2012 it had grown to 42.8 (both in thousands of dollars).
a. Use the two (year,PCPI) data points (1,38.6) and (4, 42.8) to find the linear relationship y=mx+b between x= years since 2008 and y = PCPI.
b. Interpret the slope of the line.
c. Use your linear relationship to predict PCPI in 2020.

Answer by Boreal(15235)

slope is chg y/chg x or 4.2/3 or 1.4
point slope formula y-y1=m(x-x1), m slope and (x1, y1) point
y-38.6=1.4(x-1)
y=1.4x-1.4+38.6
y=1.4x+37.2
the slope of 1.4 means the per capita icome grows $1400 a year.
2020 is 8 more years from 2012 and that makes x=12
y=16.8+37.2
=$54,000 or 54 in thousands